### v1

# Community detection in network using Szegedy quantum walk

Abstract
In a network, the vertices with similar characteristics construct communities. The vertices in a community are well-connected. Detecting the communities in a network is a challenging and important problem in the theory of complex networks. One approach to solve this problem uses the classical random walks on the graphs. In quantum computing, quantum walks are the quantum mechanical counterparts of classical random walks. In this article, we employ a variant of Szegedy’s quantum walk to develop a procedure for discovering the communities in networks. The limiting probability distribution of quantum walks assists us in determining the inclusion of a vertex in a community. We apply our procedure of community detection on a number of graphs and social networks, such as the relaxed caveman graph, $l$-partition graph, Karate club graph, dolphin’s social network, etc.
Keywords: Clustering, Community detection, Szegedy quantum walk, Complex networks and graphs.
1 Introduction
Networks are everywhere! Human beings build up networks among themselves for their progress and prosperity. They establish roads, rails, airlines, and the internet. They all form networks [1, 2]. Networks are inherent in nature. The human brain and protein-protein interaction network are two important examples of networks in nature. Due to the significance of networks, network science is an important branch of science and engineering [3]. In a network, elements of similar characteristics are connected. It distributes all the components in a network into communities. Detecting the communities in a given network is an important problem in network science [4, 5, 6, 7].
Combinatorial graphs [8, 9] form the mathematical foundation of networks. We use the words graphs and networks synonymously throughout the article. A graph is a combination of objects and connection between them, where the objects are represented by the vertices and connection are represented by edges between the vertices. We mention the mathematical definition of graphs in section 2 of this article. The real-world networks display big inhomogeneities in the distribution of edges. We observe that the distribution of edges is not only globally but also locally inhomogeneous. In a real-world network, we find high concentrations of edges within special groups of vertices and low concentrations between these groups. These groups are called communities.
Detecting the communities in a network is not easy. There are many community detection algorithms which are computationally NP-hard, or NP-complete [10, 11, 12]. The communities in graphs are often called the clusters. Therefore, the problem of community detection is also known as the graph clustering problem. Due to the existence of edges between the vertices, we prefer to use the word community instead of cluster. There are different methods of community detection discussed in the literature [13, 14, 15, 16, 17]. The random walks on graphs provide a crucial method to detect the communities [18, 19, 20, 21]. A random walker travels between the vertices following the connectivity structure of the graph. As the movement of the walker is random, at any time instance, there is a probability distribution for getting the walker at different vertices. Therefore, we get a sequence of probability distributions that converges to a limiting probability distribution under certain conditions. Usually, the inter-community edges are higher than the probability of the intra-community edges. In this way, the limiting probability distribution of random walks assists us in detecting the communities in graphs.
In quantum computing, the quantum walks [22] are the quantize versions of classical random walks on the graphs [23, 24]. There are multiple proposals of quantum walks, such as discrete-time quantum walks [25], continuous-time quantum walks [26], open quantum walks [27], etc. The Szegedy quantum walk [25], is a variant of the discrete-time quantum walk. It converts the transition matrix of a classical random walk to a unitary matrix to govern the quantum walker on graphs. In addition we can define it on all the simple graph. Therefore, the Szegedy quantum walk is an efficient way to replace the idea of a classical random walk for developing a quantum computational algorithm.
The graph clustering problem has recently been considered in the literature of quantum computing [28, 29, 30]. The quantum walks are also introduced for finding the clusters in graphs [31, 32]. Article [32] utilizes the discrete-time coined quantum walks to identify the communities in different real-world networks, such as Zachary’s karate club graph and the dolphin social networks and airline networks. The Fourier and Grover coins are used as quantum coin operator. It is demonstrated that the Fourier coin quantum walks localize within communities, which effectively reveal the modular structures of networks. In [31], the authors develop another graph clustering algorithm using discrete-time quantum random walks. Their method improves the clustering speed and accuracy. Article [30] discusses another efficient approach to clustering on directed and mixed graphs without using the idea of quantum walks.
The motivation of this work is developed from our study $G$ the Szegedy quantum walk on the Barbell graphs. A barbell graph has two communities of vertices that are joined to a vertex by two edges and a vertex. The formal description will be mentioned later in this article. We observed that the limiting probability distribution of Szegedy quantum walk from a particular initial state has higher probability values for the inter-community edges. Therefore, using the Szegedy quantum walks, we can determine the inter community edges, whose removal generates the communities in the graph. This observation assists us in developing a method for community detection in networks.
In this article, we present a new method for finding the communities in networks, which is applicable to all simple graphs with communities. There are multiple formalisms of Szegedy quantum walk, reported in literature [25]. For our constructions, we apply model discussed in [25]. To the best of our knowledge, this is the first research article where Szegedy quantum walk is applied for community detection. The quantum walk on graphs provides a limiting probability distribution on the edges of the graph. It assists us in determining a vertex to include in a community. For the numerical experiments, we apply our community detection techniques on relaxed caveman graphs, $l$-partition graphs, Zachary’s karate club graph, etc. The relaxed caveman graphs, $l$-partition graphs, are employed in the literature to examine different community detection methods [32, 5, 33]. Our result also matches with them. Due to the difference between the classical and quantum walks, we observe that the communities in Zachary’s karate club graph and the dolphin social network our communities are different from the communities derived in classical cases.
The structure of this article is as follows: The primary concepts to go through this article, such as graphs, Szegedy quantum walks, etc., are discussed in section 2. In section 3, we describe our procedures of community detection using Szegedy’s quantum walk. The section 4 contains the communities in different graphs and social networks generated by our procedure. Then we conclude the article. There are two Appendix of this article. We consider a specific quantum state depending on the graphs for fast convergence of the probability distributions to the desired limiting probability distribution. In Appendix A, we discuss the choice of the particular quantum state. In Appendix B, we explain the convergence of the probability distribution.
2 Preliminary
2.1 A brief Overview of graph theory
A graph $G=(V(G),E(G))$ is a combinatorial structure, consists of a set of vertices $V(G)$ and a set of edges $E(G)\subseteq V(G)\times V(G)$ [8]. Throughout this article, $n$ and $m$ denote the number of vertices and edges, respectively, in a graph $G$, which are the finite natural numbers. If there is a direction on an edge we describe it as a directed edge. An undirected edge has no direction on it. An undirected edge between the vertices $u$ and $v$ is denoted by $(u,v)$. A loop is an edge joining a vertex with itself. An edge with an edge-weight is called an weighted edge. A graph without any directed edge, weighted edge, and loop is called a simple graph. Pictorially, we denote a vertex by a circle which may be marked or unmarked. A line joining two vertices $u$ and $v$ denotes an undirected edge $(u,v)$. A graph is said to be a complete graph if there is an edge between any two distinct vertices. The complete graph with $n$ vertices is denoted by $K_{n}$.
In a simple graph $G$, two vertices $u$ and $v$ are said to adjacent if there is an edge $(u,v)\in E(G)$. Also, the edge $(u,v)$ is incident to the vertices $u$ and $v$. The number of incident edges to a vertex $v$ is called the degree of vertex $v$, which is denoted by $d_{G}(v)$. The neighborhood of a vertex $v$ in a simple graph $G$ is the set of all of its adjacent vertices $N(v)=\{v:(u,v)\in E(G)\}$. A path between the vertices $v_{0}$ and $v_{n}$ is a sequence of vertices and edges $v_{0},e_{1},v_{1},e_{2},v_{2},\dots e_{n},v_{n}$ where $e_{i}=(v_{i-1},v_{i})$ for $i=1,2,\dots n$, where no vertex is repeated. The number of edges in a path is called the length of a path. A graph is called connected if there is a path between any two vertices in it. In a connected graph $G$ there may be multiple paths between two vertices $u$ and $v$. The path with minimum length is called the shortest path between $u$ and $v$. A graph $G_{1}=(V(G_{1}),E(G_{1}))$ is said to be a subgraph of a graph $G=(V(G),E(G))$, if $V(G_{1})\subseteq V(G)$, and $E(G_{1})\subseteq E(G)$. A subgraph $G_{1}$ of $G$ is called an induced subgraph if $E(G_{1})=\{(u,v):(u,v)\in E(G)~\text{and}~u,v\in V(G_{1})\}$.
The directed graphs play a crucial role in this article. To denote a directed edge between the vertices $u$ and $v$ we use the symbol $\overrightarrow{(u,v)}$. A directed graph is a graph having directed edges only. We denote a directed graph by $\overrightarrow{G}=(V(\overrightarrow{G}),E(\overrightarrow{G}))$. A directed edge $\overrightarrow{(u,v)}$ is represented by an arrow from $u$ to $v$. To denote two oppositely oriented edges $\overrightarrow{(u,v)}$ and $\overrightarrow{(v,u)}$ together, we use a straight line with two opposite arrows. Given a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, the vertex $u$ is said to be the origin of $\overrightarrow{e}$, and the vertex $v$ is the terminal of the edge $\overrightarrow{e}$. We denote $u=o(\overrightarrow{e})$, and $v=t(\overrightarrow{e})$. We say that $\overrightarrow{e}$ is an incoming edge to the vertex $v$. Also, $\overrightarrow{e}$ is an outgoing edge from the vertex $u$. The indegree of a vertex $u$ is the number of incoming edges to the vertex $v$, which we denote by $d^{(i)}_{\overrightarrow{G}}(v)$. Also, the outdegree of a vertex $u$ is the number of outgoing edges from $u$, which is denoted by $d^{(o)}_{\overrightarrow{G}}(v)$. The inverse of a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, is a directed edge $\overrightarrow{e}^{-1}=\overrightarrow{(v,u)}$. Therefore, $o(\overrightarrow{e})=t(\overrightarrow{e}^{-1})$ and $t(\overrightarrow{e})=o(\overrightarrow{e}^{-1})$.
In this paper, we consider the connected, undirected graphs. A community in a graph is a set of vertices with the similar properties. In this article, we emphasize on the connectivity properties of the vertices. We assume that a vertex belongs to at most one community. Therefore, the set of communities generated by our procedure is crisp. The set of communities $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$ in a graph is a partition on $V(G)$ where $C_{i}$ are the communities for $i=1,2,\dots k$. Therefore, $\cup_{i=1}^{n}C_{i}=V(G)$ and $C_{i}\cap C_{j}=\emptyset$ for $i\neq j$. We say $\mathcal{C}$ is trivial if either $k=1$, or $C_{i}$ is a singleton set for all $i$. We identify a community $C_{i}$ with the induced subgraph $G[C_{i}]=(C_{i},E(C_{i}))$ of the graph $G$ for every community $C_{i}$, where $E(C_{i}):=\{(v,w)\in E:v,w\in C_{i}\}$. The $E(\mathcal{C})=\cup_{i=1}^{k}E(C_{i})$ is the set of intra-community edges of the graph $G$ induced by the community detection $\mathcal{C}$. The number of edges in $E(\mathcal{C})$ is denoted by $m(\mathcal{C})$. Also, $E(G)-E(C)=\{(u,v):u\in C_{i},\ v\in C_{j},\ i\neq j\}$ is the set of inter-community edges of $G$ induced by the community detection $\mathcal{C}$. [34]. The number of intra-community edges is more than the number of inter-community edges.
The community detection in graph is the task of grouping the vertices of the graph into the communities. Recall that, different procedures of community detection develops different communities of vertices in a graph. There are a number of measures to quantify the efficiency of a community detecting procedure. Recall that $\mathcal{C}$ and $m$ are the intra-community edges, and the edges in a graph $G$, respectively. The coverage is one measure to quantity, which is defined by $\operatorname{Coverage}(\mathcal{C})=\frac{m(\mathcal{C})}{m}$. The large value of $\operatorname{Coverage}(\mathcal{C})$ indicates the better quality of $\mathcal{C}$. In classical computing the graph community detection problem is an NP-hard problem when $k\geq 3$ [35].
Communities in a network depends on the connectivity properties of a network. Not all graphs preserve a structure of community in it. For examples, the complete graph with any number of vertices results a single community. Well-known graphs with community structure includes different varieties of caveman graphs, planted partition graphs, as well as different complex networks such as the Zachary’s karate club graph, etc.
2.2 An introduction to quantum walks
Let $G=(V(G),E(G))$ be a connected simple graph. We assign two opposite orientations on every undirected edge $(u,v)\in E(G)$. Therefore, from every undirected edge $(u,v)$ in $G$, we construct two directed edges $\overrightarrow{e}=\overrightarrow{(u,v)}$ and $\overrightarrow{e^{-1}}=\overrightarrow{(v,u)}$. The new directed graph is $\overrightarrow{G}$. Therefore, if $G$ has $m$ edges, then $\overrightarrow{G}$ has $2m$ directed edges. Let the edges $\overrightarrow{e_{1}},\overrightarrow{e_{2}},\dots,\overrightarrow{e_{m}},\overrightarrow{e_{1}^{-1}},\overrightarrow{e_{2}^{-1}},\dots,\overrightarrow{e_{m}^{-1}}$ are marked by the integers $0,1,2,\dots(2m-1)$. The construction of $\overrightarrow{G}$ from $G$ indicates that $d_{G}(v)=d^{(o)}_{\overrightarrow{G}}(v)=d^{(i)}_{\overrightarrow{G}}(v)$, where $d_{G}(v)$ is the degree of $v$ in $G$, as well as $d^{(o)}_{\overrightarrow{G}}(v),d^{(i)}_{\overrightarrow{G}}(v)$ are the out-degree and in-degree of $v$ in $\overrightarrow{G}$, respectively.
Let a walker is moving on the directed graph $\overrightarrow{G}$ following the graph structure. At a particular time $t$, let the walker arrives at a vertex $v$. In the next step, the walker will move to one of the adjacent vertices of $v$. Hence, the walker must follow one of the outgoing edges $\overrightarrow{e}$ from $v$. Recall that, if out-degree of $v$ is $d_{v}^{(o)}$, then there are $d_{v}^{(o)}$ out-going edges from $v$. The walker will choose one of them with probability $p(\overrightarrow{e})$. Therefore, we can define a function $p:E(\overrightarrow{G})\rightarrow(0,1]$ satisfying
$\sum_{\overrightarrow{e}:o(\overrightarrow{e})=v}p(\overrightarrow{e})=1,$ (1)
for each vertex $v\in V$. If we assume that choosing an outgoing edge from $v$ is equiprobable, then $p(\overrightarrow{e})=\frac{1}{d^{(o)}_{v}}=\frac{1}{d_{v}}$, where $v=o(\overrightarrow{e})$.
Now, we are in a position to define the Szegedy quantum walk. We represent the state of a walker by a quantum state vector. Corresponding to the directed graph $\overrightarrow{G}$ with $2m$ edges, we consider a Hilbert space $\mathcal{H}^{2m}$ of dimension $2m$. The basis of $\mathcal{H}^{2m}$ is $\left\{|0\rangle,|1\rangle,|2\rangle,\dots,|2m-1\rangle\right\}$, where
[…]
$u_{\overrightarrow{e},\overrightarrow{f}}=\begin{cases}2\sqrt{p(\overrightarrow{e})p(\overrightarrow{f^{-1}})}-\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}&\text{if}~t(\overrightarrow{f})=o(\overrightarrow{e});\\ 0,&\text{otherwise}.\end{cases}$ (3)
Here, $\delta_{(\overrightarrow{e},\overrightarrow{f})}$ is the Kronecker delta, which is defined by
$\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}=\begin{cases}1&\text{if}~\overrightarrow{e}=\overrightarrow{f};\\ 0&\text{otherwise}.\end{cases}$ (4)
[…]
In this article, the walker begins its journey with a particular initial state. Let $V_{\max}$ is the set of vertices with maximum degree in the graph $G$ and $E_{\operatorname{init}}=\{\overrightarrow{(i,j)}\mid i\in V_{\max}\}\subset E(\overrightarrow{G})$ is the set of all outgoing edges from the vertices in $V_{\max}$. The initial state is determined by
[…]
The procedure of clustering depends on the selection of the initial state. A different choice of initial state may lead us to another limiting probability distribution, which may not be useful for us. One such example is presented in Appendix A
3 Community detection on graphs
[…]
$\sum_{k=0}^{2m-1}|a_{t,k}|^{2}=1,~\text{for all}~t.$ (7)
[…]
$p_{t}(\overrightarrow{k})=\braket{\psi^{\dagger}|P_{k}|\psi}=|\langle k|\psi\rangle|^{2}=|\alpha_{t,k}|^{2}.$ (8)
[…]
$\overrightarrow{P}_{t}=(p_{t}(\overrightarrow{e_{1}}),p_{t}(\overrightarrow{e_{2}}),\dots,p_{t}(\overrightarrow{e_{n}}),p_{t}(\overrightarrow{e_{1}^{-1}}),p_{t}(\overrightarrow{e_{2}^{-1}}),\dots,p_{t}(\overrightarrow{e_{m}^{-1}})).$ (9)
Note that, every edge $e$ of the undirected graph $G$ contributed two directed edges $\overrightarrow{e}$ and $\overrightarrow{e^{-1}}$, in the directed graph $\overrightarrow{G}$. Thus, we can determine the edge probability $e$ in $G$ at $t$-th step of quantum walk as $p_{t}(e)=p_{t}(\overrightarrow{e})+p_{t}(\overrightarrow{e^{-1}})$. Therefore the edge probability vector of the undirected graph $G$ at the $t$-th step of quantum walk is represented by
$P_{t}=(p_{t}(e_{1}),p_{t}(e_{2}),\dots,p_{t}(e_{m})),$ (10)
where $p_{t}(e_{i})=p_{t}(\overrightarrow{e_{i}})+p_{t}(\overrightarrow{e_{i}^{-1}})$. Note that, due to the unitary dynamics, $P_{t}$ need not converge to a limiting distribution for $t\rightarrow\infty$ [36]. But, the time average probability distribution converges. Hence, we have the following definitions.
Definition 1
The edge probability vector at time $T$ of a graph $G$ is defined by
$\begin{split}\Pi^{(T)}&=\left(\pi_{1}^{(T)},\pi_{2}^{(T)},\dots\pi_{m}^{(T)}\right)\\ &=\frac{1}{(T+1)}\sum_{t=0}^{T}P_{t}=\frac{1}{(T+1)}\left(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\dots,\sum_{t=0}^{T}p_{t}(e_{m})\right),\end{split}$
where $\pi_{k}^{(T)}=\frac{1}{(T+1)}\sum_{t=0}^{T}p_{t}(e_{k})$ for $k=1,2,\dots m$.
Note that, the edge probability vector at time $T$ changes over time. It proceeds to a limiting distribution when $T\rightarrow\infty$. Hence, we can define the edge probability vector of a graph.
Definition 2
The edge-probability vector of a graph $G$ is defined by
$\Pi=(p(e_{1}),p(e_{2}),\dots,p(e_{m}))=\lim_{T\to\infty}\frac{1}{(T+1)}\sum_{t=0}^{T}P_{t}=\lim_{T\to\infty}\frac{1}{(T+1)}\left(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\dots,\sum_{t=0}^{T}p_{t}(e_{m})\right).$
For our computation, rather than calculating $T\to\infty$, we find $\frac{1}{T+1}\sum_{t=0}^{T}P_{t}$ for a finite $T$ to construct the limiting probability distribution. We apply the standard Euclidean distance to differentiate two edges probability vectors $\Pi^{(T)}$ and $\Pi^{(T-1)}$, which is determined by
$\begin{split}\left\lVert\Pi^{(T)}-\Pi^{(T-1)}\right\rVert&=\left\lVert\frac{1}{T+1}\sum_{t=0}^{T}P_{t}-\frac{1}{T}\sum_{t=0}^{T-1}P_{t}\right\rVert\\[6.0pt] &=\sqrt{(\pi_{1}^{(T)}-\pi_{1}^{(T-1)})^{2}+(\pi_{2}^{(T)}-\pi_{2}^{(T-1)})^{2}+\cdots+(\pi_{m}^{(T)}-\pi_{m}^{(T-1)})^{2}}.\end{split}$ (11)
To construct the edge probability vector of a graph $G$, we compute a sequence of edge probability vectors $\{\Pi^{(t)}\}_{t=0}^{T}$ for time $t$, where
$\left|\Pi^{(T)}-\Pi^{(T-1)}\right|<\epsilon.$ (12)
Here, we consider $\epsilon$ as a numerical error which is usually a very small number. For our numerical experiments, we consider $\epsilon=0.0001$. The consideration of numerical error assists us to find the accurate value of $\Pi$ to three decimal places. In 1(b), we present the edge-probabilities of the graph in 1(a). The convergence of the probability vectors are discussed in Appendix B.
(a) Graph $G$
(b) Edge probability in the limiting distribution $\Pi$ is depicted as bar diagram.
(c) The vertex $5$ in $G$ has the maximum degree. The green vertices are in $\operatorname{Nbd}(5)$, which is generated following Procedure 1.
(d) The path weights of the paths between different vertices in $G$ are calculated by equation (13). They are plotted using bar diagram.
(e) Communities in $G$ are extracted without refinement. Communities are represented by three different colors.
(f) Communities in $G$ are extracted with refinement.Communities are represented by three different colors.
Figure 1: Different steps for generating communities on the graph $G$, depicted in Sub-figure 1(a), are presented.
Now, we include a number graph theoretic ideas in construction of our communities. A vertex with degree $1$ is adjacent to only one vertex, which may belong to a community. A vertex with degree two lies on a path between two neighboring vertices. These three vertices may form a single community, or the neighboring vertices may belong to two different communities. But, a vertex with greater degree is adjacent to many other vertices. Thus, it belong to only one community. It insists us to arrange the vertices in a descending order of their degrees. In every step, we consider the highest degree vertex among the vertices which are not allotted to a community and generate a community for it.
In addition, the above paragraph justifies that the community containing a vertex with highest degree also contains all its neighboring vertices. But, there may be a number of vertices in a community, which are not an immediate neighbor of the vertex with highest degree in it. Also, there are non-neighboring vertices which are not in the community. When a vertex is in the community we observe that most of its neighbours are also in the same community. Therefore, we need a condition for including a given vertex in a particular community. Recall from subsection 2.1 that $N(v_{i})=\{v:(u,v)\in E(G)\}$. Let $\operatorname{Nbd}(v_{i})$ be the neighborhood of $v_{i}$ in the community of vertices containing $v_{i}$. If $|N(v_{j})\cap\operatorname{Nbd}(v_{i})|\geq\frac{|N(v_{j})|}{2}$ holds, we assume that $v_{j}$ is in the community of $v_{i}$.
We make the final selection of a vertex in a community based on the edge probability vector $\Pi$ generated by quantum walk. We determine the shortest path between $v_{i}$ and $v_{j}$. A path consists of a number of incident edges. The walker follows the path means it follows all the subsequent edges. As there is no non-Markovian componenet in our quantum walk, the probability of getting the walker on the edges of a path is independent to each other. Therefore, the probability of getting the walker on a path is the product of the probability of getting the walker on the edges. It leads us to define the path weight between two vertices.
Definition 3
Let the shortest path between $u$ and $v$ in graph $G$ be $u=u_{0},e_{1},u_{1},e_{2},u_{2},\dots e_{n},u_{n}=v$. We define the path weight between the vertices $u$ and $v$ by
$PW(u,v)=\frac{1}{d_{G}(v)}\prod_{i=1}^{n}p(e_{i}),$ (13)
where $p(e_{i})$ is the probability corresponding to the edge $e_{i}$ in the edge probability vector $\Pi$.
Note that, as all $0<p(e_{i})<1$ and $d_{G}(v)\geq 1$ we have $0<PW(u,v)<1$ for all pairs of vertices $u$ and $v$. The explanation for dividing $\frac{1}{d_{G}(v)}$ will be discussed after Procedure Procedure 1.
Now, we are in a position to determine the communities in a network. Given a vertex $v_{i}$ the community consists of $v_{i}$ is denoted by $C(v_{i})$. Initially, $C(v_{i})=\{v_{i}\}$. Then, $C(v_{i})$ grows maintaining some conditions. We choose the vertices $v_{i}$ according to their degrees. Our method is described in the procedure mentioned below.
Procedure 1
Recall that $N(v)$ denotes the subset of vertices which are adjacent to $v$ in $G$. The procedure for detecting the communities consists of the following steps:
  1. 1.
    Arrange the vertices of the graph $G$ in the descending order of the vertex degrees. Therefore, the vertex set of $G$ is $V=\{v_{1},v_{2},\dots v_{n}\}$, where $d_{G}(v_{i})\geq d_{G}(v_{i+1})$ for $i=1,2,\dots(n-1)$.
  2. 2.
    Repeat the following steps if $V\neq\emptyset$.
    1. (a)
      Consider the first element of $V$, say $v_{i}$.
    2. (b)
      Construct the set of vertices $\operatorname{Nbd}(v_{i})=N(v_{i})$.
    3. (c)
      Let $v_{j}\notin N(v_{i})$ is a vertex satisfying
      $|N(v_{j})\cap\operatorname{Nbd}(v_{i})|\geq\frac{|N(v_{j})|}{2},$ (14)
      then add $v_{j}$ in $\operatorname{Nbd}(v_{i})$.
    4. (d)
      Construct the community with $v_{i}$, by setting $C(v_{i})=\{v_{i}\}$.
    5. (e)
      Add all $v_{j}\in\operatorname{Nbd}(v_{i})$ in $C(v_{i})$ if $PW(v_{i},v_{j})<q$.
    6. (f)
      Update the set $V$ by removing all the vertices in $C(v_{i})$.
  3. 3.
    Return the communities of $C(v_{i})$ which are formed in Step 2.
It is clear from the Procedure 1 mentioned above that we construct a set of vertices $\operatorname{Nbd}(v_{i})$ based on equation (14), which assists us to construct the community $C(v_{i})$. The quantum walk is useful to calculate the probability values in finding $PW(v_{i},v_{j})$ mentioned in equation (13).
In the above procedure $q$ is a parameter. If $q=0$ then the number of communities is equal to the number of vertices in the graph, that is every community consists of single vertices. When $q>\max_{(u,v)\in E(G)}\{PW(u,v)\}$ then $C(v_{i})=\operatorname{Nbd}(v_{i})$.
In equation (13), we multiply $\frac{1}{d_{G}(v)}$ to normalize node influence. It provides more preference to a high-degree node to be added in $C(v_{i})$, because a node with higher degree reduces the product $\prod_{i=1}^{n}p(e_{i})$ than a vertex with lower degree.
After constructing a community $C(v_{i})$ using Procedure Procedure 1, it may be observed that a number of additional vertices are in the community. For example, consider Figure 1. Here, the vertex $6$ is added in the community $C(1)$. But the structure of the graph suggests that vertex $6$ should form a separate community. Therefore, need to remove it from $C(1)$. To discard those vertices we perform a refinement using the below procedure.
Procedure 2
Let a community $C(v_{i})$ is created using Step $1$ and $2$ of Procedure 1. Apply the following step before proceeding to step $3$.
  1. 1.
    For any vertex $v_{j}\in C(v_{i})$ if $d_{C(v_{i})}(v_{j})\leq\frac{d_{G}(v_{j})}{2}$ then remove $v_{j}$ from $C(v_{i})$. Here, $C(v_{i})$ denotes the induced subgraph of $G$, generated by $C(v_{i})$.
Therefore, our method of community detection is a combination of the following steps, mentioned in the below procedure.
Procedure 3
Let $G$ be a simple undirected graph with $n$ vertices and $m$ edges. To find the communities in $G$ perform the following steps:
  1. 1.
    Generation of a directed graph: Convert the graph $G$ to a directed graph $\overrightarrow{G}$ by assigning opposite orientations on every undirected edge.
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
    Generation of limiting probability distribution: For a fixed error term $\epsilon$, we apply equation (12) to stop quantum walk and generate the limiting probability distribution $\Pi$ on the edges mentioned in Definition 2.
  5. 5.
    Generation of communities Apply Procedure 1 and 2 for generating communities.
4 Numerical experiments
We have already mentioned that not all graphs admits community structure in it. We consider a number of graphs with a community structure. We also consider a graph having no trivial community structure. We initiate our discussion with a few standard examples. Some of the graphs considered in this article are the random graphs. Due to their generation procedure, they may have directed edges or loops. But, our procedure is applicable for undirected simple graphs. During the network generation process, we remove the loops and directed edges if exist.
For every graph we include a figure with at least four sub-figures. Below, we describe them:
  1. 1.
    In Sub-figure (a), we draw the example graph, which we consider as our input.
  2. 2.
    In Sub-figure (b), we include a bar diagram representing the limiting edge-probability distribution $\Pi$, mentioned in Definition 2 after completion of quantum walk.
  3. 3.
    In Sub-figure (c), we draw a bar diagram mentioning the path weights between selected vertices in the $\operatorname{Nbd}$ of the community generating vertex. The equation of path probability is mentioned in equation (13).
  4. 4.
    In Sub-figure (d), we draw the graph with communities in it. The vertices in same community have same color. There are communities with single vertices. We mark all of them in gray.
We use NetworkX, a Python library, for all the calculations related to graphs [37].
4.1 Barbell graphs
A barbell graph [38] is generated by joining $r$ copies of a complete graph $K_{s}$ with a particular vertex, called central vertex. We join the central vertex to the $r$ vertices belonging to different $K_{s}$ by $r$ edges. For example in 2(a), we depict a barbell graph, that has two complete graphs each of five vertices. We add two vertices from each of the $K_{5}$ graphs with a vertex $5$ by two edges. Here vertex $5$ acts as the central vertex. Therefore, the barbell graph under our consideration has eleven vertices and twenty-two edges.
Considering $q=\frac{1}{m}=0.0455$ in Procedure 3, we obtain three distinct communities in the barbel graph network. Two distinct communities arise from two complete graphs.The central vertex forms a community with a single vertex. The coverage values corresponding to $q=0.0455$ is $0.909$. In 2(d) we present the communities generated by Procedure 3.
(a) A barbell graph where two complete graph $k_{5}$ are joined to a central vertex $5$ by two edge $(5,6),(5,4)$ .
(b) Edge probabilities in the limiting distribution $\Pi$ are depicted as bar diagram for the graph in Subfigure 2(a).
(c) Path weight between different vertices of the for generating the communities from the graph in 2(a).
(d) Communities in the graph of Subfigure 2(a) are extracted with three different colors.
Figure 2: Community detection on a barbell graph
4.2 Relaxed caveman graph
The idea of caveman graphs arises from the social networks [39, 40]. The connected caveman graph is generated by creating $k$ complete subgraphs with equal number of vertices in each. Then a single edge in each clique is rewired to a node in an adjacent clique. These cliques are called caves. The relaxed caveman graph is a variant of the connected caveman graphs. The network is composed of $k$ nodes organized into $k$ caves. For every edge there is a probability $p$ of rewiring, that is removing an edge and adding it to any vertex keeping one end point fixed. A relaxed caveman graph is depicted in 3(a).
The example graph, depicted in 3(a), has $15$ vertices and $30$ edges. The vertices are distributed into $3$ caves. Every cave contains $5$ vertices in it. We consider that the rewiring probability is $0.2$. The example graph has $4$ inter-community edges.
Note that, in this case $\frac{1}{m}=0.033$. The communities are marked with three different colors in Sub-figure 3. Decreasing the value of $q_{2}$ to $0.0166$ we find more communities in our network. Coverage are $0.866$ and $0.766$ for $q=0.0333$ and $0.0166$ respectively.
(a) Relaxed caveman graph having three caves with rewiring probability $p=0.2$.
(b) Edge probability in the limiting distribution $\Pi$ is depicted as bar diagram for the graph in Subfigure 3(a).
(c) Path weight between different vertices of the for generating the communities from the graph in 3(a).
(d) Communities in the graph of Subfigure 3(a) are extracted with three different colors.
(e) Considering $q=\frac{1}{2m}$, the path weight between different vertices of the graph for generating the communities from the graph in 3(a).
(f) Communities in the graph of Subfigure 3(a) are extracted with four different colors.
Figure 3: Community detection in a relaxed caveman graph.
4.3 Planted $l$-partition graph
The planted partition models of networks offer another controllable artificial input for evaluating and benchmark the community detection algorithms. The Erdős-Rényi model of random graphs [41] with $n$ vertices is generated by choosing each of the $\binom{n}{2}$ possible edges independently and randomly to be included in the graph with probability $p$. The planted $l$-partition model is a modified version of the Erdős–Rényi model, which is specifically designed to generate communities in graph. To generate $l$ communities in a graph with $k$ vertices in each, we need $n=l\times k$ vertices. Now, we fix two probabilities $p$ and $q$ with $p>q$. Existence of an edge between two vertices belonging to same cluster is determined by probability $p$. An edge between two vertices from two different cluster occurs in probability $q$. A planted $l$-partition graph is depicted in 4(a).
We consider a sample of a planted partition graph, which in 4(a). It has twelve vertices and fourteen edges. Also, values of $p_{1}$ and $p_{2}$ are $0.4$ and $0.1$, respectively. The vertices are divided into three subset each having four vertices in it.
The vertices with the highest degrees in the graph are identified as 0, and $7$. Among them, vertex 0 is randomly selected as the first vertex to initiate Procedure 1 for constructing a community. A bar diagram is employed to illustrate the route probability distribution. Based on the graphs structural connectivity, a distinct community can be obtained for $q>0.071$. The coverage values corresponding to $q=0.071$ is $0.857$.
(a) A planted partition graph with $p_{1}=0.4$ and $p_{2}=0.1$.
(b) Edge probability in the limiting distribution $\Pi$ is depicted as bar diagram for the graph in Subfigure 4(a).
(c) Path weight between different vertices of the for generating the communities from the graph in 4(a).
(d) Communities in the graph of Subfigure 4(a) are extracted with three different colors.
Figure 4: Community detection in a planted partition graph.
4.4 Zachary’s karate club graph
The Zachary’s karate-club network [42] is another well-studied graph considered in the literature of community detection [33, 32]. This is a social network of university karate club members. Each of the $34$ vertices represents a club member. There are $78$ edges represent the relations between the members. Article [33] presents three communities in the Zachary’s karate-club network. In contrast, article [32] utilize a different community detection technique to analyze the same network that found two communities only.
Considering $q=\frac{1}{m}=0.0128$ in Procedure 3, we obtain three distinct communities in the Zachary’s karate-club network. Two communities have fourteen and eighteen vertices, and one community consists of a single vertex. The coverage values corresponding to $q=0.0128$ is found to be $0.859$. In 5(d) we present the communities determined by Procedure 3. In contrast, the article [33] determines three communities having five, ten, and nineteen vertices. Furthermore, article [32] develops two communities consist of sixteen and eighteen vertices.
(a) Zachary’s karate club network.
(b) The bar diagram for the graph in Subfigure 5(a) represents the edge probability in the limiting distribution $\Pi$.
(c) In 5(a), the path weight between various vertices of the graph is used to generate the communities.
(d) Three distinct colors are used to extract communities from the graph in Subfigure 5(a).
Figure 5: Karate club graph are considered to generate communities.
4.5 Les Misérable graph
6(a) represents the “Les Misérable” graph. This network was demonstrated in [43] considering the characters in Victor Hugo’s novel “Les Misérable” as vertices. Any pair of characters that occur in the same chapter of the novel are connected by an edge. This graph has seventy seven vertices and two hundred fifty four edges.
Article [33] considers the Les Misérable graph for community detection. It determines six communities in the graph. We use Procedure 3 on the Les Misérable graph considering $q=\frac{1}{m}=0.0039$ for determining the communities. It generates eighteen communities in the graph. Among them there are three communities with vertices eight, twelve, and forty-four. Remaining fifteen communities have single vertices in each. In 6(d), we present the communities in the Les Misérable graph determined by our procedure. The coverage values corresponding to $q=0.0039$ is found to be $0.752$.
(a) Les Misérable graph
(b) The graph in Subfigure 6(a) has a bar diagram that shows the edge probability in the limiting distribution $\Pi$.
(c) In 6(a), the path weight between various vertices of the graph is used to generate the communities.
(d) The three communities are represented by three distinct colors in the graph shown in Subfigure 6(a), while all other single-node clusters are shown by gray.
Figure 6: Les Misérable graph are utilized to generate communities.
4.6 Dolphin’s social network
7(a) illustrates the dolphin’s social network which represents a social network among sixty two bottlenose dolphins, which are considered as nodes. Each of the one hundred fifty nine edges signifies a frequent social interaction between two dolphins spotted together more often than by chance [44].
Article [45] determines four communities in the dolphin’s social network. These communities consists of six, thirteen, twenty two and twenty one vertices. Considering $q=\frac{1}{m}=0.0063$ in Procedure 3, we obtain five distinct communities in dolphin social network. There are three communities having twelve, twenty-one, and twenty-seven vertices, and two communities consists of a single vertex in each. These single-vertex communities exist because the respective vertices are associated with several communities and hence cannot be unequivocally allocated to any one of the three communities. The coverage values corresponding to $q=0.0063$ is $0.824$. 7(a) presents the communities generated by our procedure.
(a) Dolphin social network
(b) The bar diagram for the graph in Subfigure 7(a) represents the edge probability in the limiting distribution $\Pi$.
(c) In 7(a), the path weight between various vertices of the graph is used to generate the communities.
(d) The three communities are represented by three distinct colors in the graph shown in Subfigure 7(a), while the other two single-node communities are shown by gray.
Figure 7: Dolphin Network are utilized to generate communities.
4.7 Euroroad network
We consider Euroroad data from [46] for our experiment. The road interactions are represented by the vertices in the network. Every edge denotes a road that links two junctions. Our road network contains $1174$ nodes and $1417$ edges, which is not connected. Hence, we study the largest connected component in this network containing $1039$ vertices and $1305$ edges. This component is depicted in 8(a).
Applying Procedure 3, we find three hundred eighty-nine communities. The largest community consists of thirty-five vertices. The coverage values corresponding to $q=0.00077$ is found to be $0.540$.
(a) Connected Road network
(b) The bar diagram for the graph in Subfigure 8(a) represents the edge probability in the limiting distribution $\Pi$.
(c) In 8(a), the path weight between various vertices of the graph is used to generate the communities.
(d) All communities are represented by distinct colors in the graph shown in Subfigure 8(a), while the single-node communities are shown by gray.
Figure 8: Road Network are utilized to generate communities.
5 Conclusion
In this article, we introduce a new procedure that utilizes Szegedy quantum walks to detect the communities in a network. There are multiple models of quantum walks in the literature. We employ the Szegedy quantum walks on graphs for our investigation, as it is defined on all graphs.
The limiting probability distribution of the Szegedy quantum walks assists us in finding the communities. This distribution depends on the choice of the initial state of the quantum walker. We assume that the walker begins its journey from the vertices with the highest degree. Therefore, the initial quantum state as well as the limiting probability distribution depends on the structure of the graph. The limiting probability distribution provides the edge probabilities for each edge in the graph. We use it in Procedure 3 to extract the communities from the graph. This procedure is a combination of two subroutines, which are Procedure 1 and Procedure 2. In Procedure 1, we find the communities using quantum walks. Then, Procedure 2 refines the preliminary communities. We apply Procedure 3 on different graphs to identify the communities in them.
The problem of community detection is an important problem in the study of networks. The communities assists us to understand the structure of the networks. In the future, the interested reader may apply our community detection procedure in the investigation of different complex networks.
Acknowledgment
MSR is thankful to Ministry of Education, Government of India for a doctoral fellowship.
Data availability statement
No dataset if generated during this work. All the programs to plot the figures are available in GitHub https://github.com/dosupriyo/Community-detection-using-Szegedy-quantum-walk.git.
Appendix A Dependency on the initial state
0123456
(a) An example graph
[…]
(0, 1) 0.1274 0.1625
(0, 2) 0.1274 0.1625
(1, 3) 0.1274 0.1131
(1, 4) 0.1274 0.1131
(2, 5) 0.1274 0.1131
(2, 6) 0.1274 0.1131
(3, 4) 0.1274 0.1116
(5, 6) 0.1274 0.1116
(b) Limiting probability distributions on the graph in 9(a) for different choice of initial state.
Figure 9: We calculate limiting probability distributions mentioned in Definition 2 for a graph in 9(a) with different choice of initial states, which are mentioned in equation (5) and (15). We receive two different probability distributions in two different cases. They are in the table 9(b).
In this article, we consider a particular initial state, which is mentioned in equation (5). But, we had many other options of initial state of the walker. For example the equal superposition state as follows:
[…]
Equation (15) indicates that the walker will start from all the edges selected uniformly.
[…]
Appendix B Convergence of probability distributions
In Definition 2, we claim that the edge probability vectors, mentioned in Definition 1 will converge to a vector over time. The Markovian nature of the considered quantum walk is the key reason for this convergence. We present this incident via a number of diagrams. depicted in Figure 10.
We consider six network for our investigations. From every graph we select a random sample of four edges. For different time instance every edge has an edge probability value. We observe that all these probability values converges over time for any particular edge. Find Figure 10 for the detail descriptions.
(a) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the barbell graph, depicted in 2(a). We observe the convergence from $T=743$.
(b) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the relaxed caveman graph, depicted in 3(a). We observe the convergence from $T=1329$.
(c) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the planted partition graph, depicted in 4(a). We observe the convergence from $T=1094$.
(d) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the karate club graph, depicted in 5(a). We observe the convergence from $T=882$.
(e) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the Les Misérable graph, depicted in 6(a). We observe the convergence from $T=714$.
(f) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the dolphins Graph, depicted in 7(a). We observe the convergence from $T=705$.
(g) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the road network, depicted in 8(a). We observe the convergence from $T=322$.
Figure 10: Convergence of edge probability distributions, which is mentioned in Definition 2.

### v2

@@
 Detecting the communities in a network is a challenging and important problem in the theory of complex networks.
-One approach to solve this problem uses the classical random walks on the graphs.
+One approach to solving this problem uses the classical random walks on graphs.
 In quantum computing, quantum walks are the quantum mechanical counterparts of classical random walks.
@@
 The limiting probability distribution of quantum walks assists us in determining the inclusion of a vertex in a community.
-We apply our procedure of community detection on a number of graphs and social networks, such as the relaxed caveman graph, $l$-partition graph, Karate club graph, dolphin’s social network, etc.
-Keywords: Clustering, Community detection, Szegedy quantum walk, Complex networks and graphs.
+We apply our community detection procedure to a variety of graphs and social networks, including the relaxed caveman graph, $l$-partition graph, Karate club graph, and the dolphin’s social network, among others.
+Keywords: Clustering, Community detection, Szegedy quantum walk, Complex networks.
 1 Introduction
@@
 They establish roads, rails, airlines, and the internet.
-They all form networks [1, 2].
+They all are different networks [1, 2].
 Networks are inherent in nature.
-The human brain and protein-protein interaction network are two important examples of networks in nature.
-Due to the significance of networks, network science is an important branch of science and engineering [3].
-In a network, elements of similar characteristics are connected.
-It distributes all the components in a network into communities.
+The human brain and protein-protein interaction network are two important examples of networks in nature [3].
+In a network, elements of similar characteristics are well-connected.
+It distributes all the elements in a network into communities.
 Detecting the communities in a given network is an important problem in network science [4, 5, 6, 7].
@@
 We use the words graphs and networks synonymously throughout the article.
-A graph is a combination of objects and connection between them, where the objects are represented by the vertices and connection are represented by edges between the vertices.
-We mention the mathematical definition of graphs in section 2 of this article.
+A graph is a combination of objects and connection between them, where the objects are represented by the vertices and connection are represented by edges between them.
 The real-world networks display big inhomogeneities in the distribution of edges.
-We observe that the distribution of edges is not only globally but also locally inhomogeneous.
-In a real-world network, we find high concentrations of edges within special groups of vertices and low concentrations between these groups.
-These groups are called communities.
-Detecting the communities in a network is not easy.
-There are many community detection algorithms which are computationally NP-hard, or NP-complete [10, 11, 12].
-The communities in graphs are often called the clusters.
-Therefore, the problem of community detection is also known as the graph clustering problem.
-Due to the existence of edges between the vertices, we prefer to use the word community instead of cluster.
-There are different methods of community detection discussed in the literature [13, 14, 15, 16, 17].
+The concentrations of edges within special groups of vertices is high.
+We call them communities, or clusters.
+Due to the existence of connections between the vertices, we prefer to use the word community instead of cluster.
+Detecting the communities in a network is not an easy task.
+There are different methods of community detection discussed in the literature [10, 11, 12, 13, 14, 15].
+Many of the community detection algorithms are computationally NP-hard, or NP-complete [16, 17, 18].
 The random walks on graphs provide a crucial method to detect the communities [19, 20, 21, 22].
 A random walker travels between the vertices following the connectivity structure of the graph.
-As the movement of the walker is random, at any time instance, there is a probability distribution for getting the walker at different vertices.
+As the movement of the walker is random, at any time-instance, there is a probability distribution for getting the walker at different vertices.
 Therefore, we get a sequence of probability distributions that converges to a limiting probability distribution under certain conditions.
 Usually, the inter-community edges are higher than the probability of the intra-community edges.
-In this way, the limiting probability distribution of random walks assists us in detecting the communities in graphs.
+In this way, the limiting probability distribution of random walks assists us in detecting the communities.
 In quantum computing, the quantum walks [23] are the quantize versions of classical random walks on the graphs [24, 25].
-There are multiple proposals of quantum walks, such as discrete-time quantum walks [25], continuous-time quantum walks [26], open quantum walks [27], etc.
-The Szegedy quantum walk [25], is a variant of the discrete-time quantum walk.
-It converts the transition matrix of a classical random walk to a unitary matrix to govern the quantum walker on graphs.
+The quantum walks are applied in quantum communication [26, 27], and computation [28, 29, 30].
+There are multiple proposals of quantum walks, such as discrete-time quantum walks [31], continuous-time quantum walks [32], open quantum walks [33, 34], etc.
+The Szegedy quantum walks [31] converts the transition matrix of a classical random walk to a unitary matrix to govern the quantum walker on graphs.
 In addition we can define it on all the simple graph.
-Therefore, the Szegedy quantum walk is an efficient way to replace the idea of a classical random walk for developing a quantum computational algorithm.
+Therefore, the Szegedy quantum walk is an efficient way to replace the idea of random walk for developing a quantum algorithm.
 The graph clustering problem has recently been considered in the literature of quantum computing [35, 36, 37].
 The quantum walks are also introduced for finding the clusters in graphs [38, 39].
-Article [32] utilizes the discrete-time coined quantum walks to identify the communities in different real-world networks, such as Zachary’s karate club graph and the dolphin social networks and airline networks.
-The Fourier and Grover coins are used as quantum coin operator.
+Article [39] utilizes the discrete-time coined quantum walks to identify the communities in different real-world networks.
+The Fourier and Grover coins are used as the quantum coin operators.
 It is demonstrated that the Fourier coin quantum walks localize within communities, which effectively reveal the modular structures of networks.
-In [31], the authors develop another graph clustering algorithm using discrete-time quantum random walks.
-Their method improves the clustering speed and accuracy.
-Article [30] discusses another efficient approach to clustering on directed and mixed graphs without using the idea of quantum walks.
-The motivation of this work is developed from our study $G$ the Szegedy quantum walk on the Barbell graphs.
-A barbell graph has two communities of vertices that are joined to a vertex by two edges and a vertex.
-The formal description will be mentioned later in this article.
+Article [38] improves the clustering speed and accuracy.
+Article [37] discusses another efficient approach to clustering on directed and mixed graphs without using quantum walks.
+The motivation of this work comes from the Szegedy quantum walk on the Barbell graphs.
+A barbell graph has two communities of vertices that are joined to a central vertex by two inter-community edges.
+Ref.
+2(a).
 We observed that the limiting probability distribution of Szegedy quantum walk from a particular initial state has higher probability values for the inter-community edges.
-Therefore, using the Szegedy quantum walks, we can determine the inter community edges, whose removal generates the communities in the graph.
+Their removal generates the communities in the graph.
 This observation assists us in developing a method for community detection in networks.
 In this article, we present a new method for finding the communities in networks, which is applicable to all simple graphs with communities.
-There are multiple formalisms of Szegedy quantum walk, reported in literature [25].
-For our constructions, we apply model discussed in [25].
+For our constructions, we apply the model of Szegedy quantum walks discussed in [31].
 To the best of our knowledge, this is the first research article where Szegedy quantum walk is applied for community detection.
@@
 It assists us in determining a vertex to include in a community.
-For the numerical experiments, we apply our community detection techniques on relaxed caveman graphs, $l$-partition graphs, Zachary’s karate club graph, etc.
-The relaxed caveman graphs, $l$-partition graphs, are employed in the literature to examine different community detection methods [32, 5, 33].
-Our result also matches with them.
-Due to the difference between the classical and quantum walks, we observe that the communities in Zachary’s karate club graph and the dolphin social network our communities are different from the communities derived in classical cases.
-The structure of this article is as follows: The primary concepts to go through this article, such as graphs, Szegedy quantum walks, etc., are discussed in section 2.
-In section 3, we describe our procedures of community detection using Szegedy’s quantum walk.
-The section 4 contains the communities in different graphs and social networks generated by our procedure.
+For the numerical experiments, we apply our community detection techniques on different networks which are already used in literature [39, 5, 40].
+Our results also matches with them.
+Due to the difference between the random walks and quantum walks, we observe that the communities are different in some cases.
+The structure of this article is as follows: The primary concepts are discussed in section 2.
+In section 3, we describe our procedures of community detection.
+The section 4 contains the communities in different networks generated by our procedure.
 Then we conclude the article.
-There are two Appendix of this article.
-We consider a specific quantum state depending on the graphs for fast convergence of the probability distributions to the desired limiting probability distribution.
-In Appendix A, we discuss the choice of the particular quantum state.
-In Appendix B, we explain the convergence of the probability distribution.
+In Appendix we discuss the choice of our initial quantum state, and the convergence of the probability distribution.
 2 Preliminary
 2.1 A brief Overview of graph theory
-A graph $G=(V(G),E(G))$ is a combinatorial structure, consists of a set of vertices $V(G)$ and a set of edges $E(G)\subseteq V(G)\times V(G)$ [8].
-Throughout this article, $n$ and $m$ denote the number of vertices and edges, respectively, in a graph $G$, which are the finite natural numbers.
-If there is a direction on an edge we describe it as a directed edge.
-An undirected edge has no direction on it.
+A graph $G=(V(G),E(G))$ consists of a set of vertices $V(G)$ and a set of edges $E(G)\subseteq V(G)\times V(G)$ [8].
+Throughout this article, $n$ and $m$ denote the number of vertices and edges, respectively.
 An undirected edge between the vertices $u$ and $v$ is denoted by $(u,v)$.
 A loop is an edge joining a vertex with itself.
-An edge with an edge-weight is called an weighted edge.
-A graph without any directed edge, weighted edge, and loop is called a simple graph.
-Pictorially, we denote a vertex by a circle which may be marked or unmarked.
-A line joining two vertices $u$ and $v$ denotes an undirected edge $(u,v)$.
-A graph is said to be a complete graph if there is an edge between any two distinct vertices.
-The complete graph with $n$ vertices is denoted by $K_{n}$.
-In a simple graph $G$, two vertices $u$ and $v$ are said to adjacent if there is an edge $(u,v)\in E(G)$.
+A graph with the undirected edges, and without loops is called a simple graph.
+We do not consider weighted graphs, in this article.
+A graph with $n$ vertices is said to be a complete graph, $K_{n}$ if there is an edge between any two distinct vertices.
+In a simple graph $G$, two vertices $u$ and $v$ are said to be adjacent if there is an edge $(u,v)$.
 Also, the edge $(u,v)$ is incident to the vertices $u$ and $v$.
-The number of incident edges to a vertex $v$ is called the degree of vertex $v$, which is denoted by $d_{G}(v)$.
-The neighborhood of a vertex $v$ in a simple graph $G$ is the set of all of its adjacent vertices $N(v)=\{v:(u,v)\in E(G)\}$.
+The number of incident edges to a vertex $v$ is called the degree of $v$, which is denoted by $d_{G}(v)$.
+The neighborhood of a vertex $v$ in $G$ is the set $N(v)=\{v:(u,v)\in E(G)\}$.
 A path between the vertices $v_{0}$ and $v_{n}$ is a sequence of vertices and edges $v_{0},e_{1},v_{1},e_{2},v_{2},\dots e_{n},v_{n}$ where $e_{i}=(v_{i-1},v_{i})$ for $i=1,2,\dots n$, where no vertex is repeated.
 The number of edges in a path is called the length of a path.
-A graph is called connected if there is a path between any two vertices in it.
-In a connected graph $G$ there may be multiple paths between two vertices $u$ and $v$.
-The path with minimum length is called the shortest path between $u$ and $v$.
+A graph is called connected if there is a path between any two vertices.
+The path between $u$ and $v$ with minimum length is called the shortest path.
 A graph $G_{1}=(V(G_{1}),E(G_{1}))$ is said to be a subgraph of a graph $G=(V(G),E(G))$, if $V(G_{1})\subseteq V(G)$, and $E(G_{1})\subseteq E(G)$.
 A subgraph $G_{1}$ of $G$ is called an induced subgraph if $E(G_{1})=\{(u,v):(u,v)\in E(G)~\text{and}~u,v\in V(G_{1})\}$.
-The directed graphs play a crucial role in this article.
-To denote a directed edge between the vertices $u$ and $v$ we use the symbol $\overrightarrow{(u,v)}$.
-A directed graph is a graph having directed edges only.
-We denote a directed graph by $\overrightarrow{G}=(V(\overrightarrow{G}),E(\overrightarrow{G}))$.
-A directed edge $\overrightarrow{(u,v)}$ is represented by an arrow from $u$ to $v$.
-To denote two oppositely oriented edges $\overrightarrow{(u,v)}$ and $\overrightarrow{(v,u)}$ together, we use a straight line with two opposite arrows.
-Given a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, the vertex $u$ is said to be the origin of $\overrightarrow{e}$, and the vertex $v$ is the terminal of the edge $\overrightarrow{e}$.
-We denote $u=o(\overrightarrow{e})$, and $v=t(\overrightarrow{e})$.
-We say that $\overrightarrow{e}$ is an incoming edge to the vertex $v$.
-Also, $\overrightarrow{e}$ is an outgoing edge from the vertex $u$.
-The indegree of a vertex $u$ is the number of incoming edges to the vertex $v$, which we denote by $d^{(i)}_{\overrightarrow{G}}(v)$.
-Also, the outdegree of a vertex $u$ is the number of outgoing edges from $u$, which is denoted by $d^{(o)}_{\overrightarrow{G}}(v)$.
-The inverse of a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, is a directed edge $\overrightarrow{e}^{-1}=\overrightarrow{(v,u)}$.
+A directed edge from $u$ to $v$ is denoted by $\overrightarrow{(u,v)}$.
+All the edges in a directed graph $\overrightarrow{G}=(V(\overrightarrow{G}),E(\overrightarrow{G}))$ is directed.
+Given a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, the vertices $u=o(\overrightarrow{e})$, and $v=t(\overrightarrow{e})$ are said to be the origin and the terminal of $\overrightarrow{e}$, respectively.
+The indegree of a vertex $v$ is denoted by $d^{(i)}_{\overrightarrow{G}}(v)$, which is the number of edges $\overrightarrow{e}$ such that $v=t(\overrightarrow{e})$.
+Also, the outdegree of a vertex $v$ is denoted by $d^{(o)}_{\overrightarrow{G}}(v)$, which is the number of edges $\overrightarrow{e}$ such that $v=o(\overrightarrow{e})$.
+The inverse of a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, is $\overrightarrow{e}^{-1}=\overrightarrow{(v,u)}$.
 Therefore, $o(\overrightarrow{e})=t(\overrightarrow{e}^{-1})$ and $t(\overrightarrow{e})=o(\overrightarrow{e}^{-1})$.
-In this paper, we consider the connected, undirected graphs.
-A community in a graph is a set of vertices with the similar properties.
-In this article, we emphasize on the connectivity properties of the vertices.
-We assume that a vertex belongs to at most one community.
-Therefore, the set of communities generated by our procedure is crisp.
-The set of communities $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$ in a graph is a partition on $V(G)$ where $C_{i}$ are the communities for $i=1,2,\dots k$.
-Therefore, $\cup_{i=1}^{n}C_{i}=V(G)$ and $C_{i}\cap C_{j}=\emptyset$ for $i\neq j$.
-We say $\mathcal{C}$ is trivial if either $k=1$, or $C_{i}$ is a singleton set for all $i$.
+In this article, we assume that a vertex belongs to at most one community.
+The set of communities $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$ in a graph is a partition on $V(G)$, such that, $\cup_{i=1}^{n}C_{i}=V(G)$ and $C_{i}\cap C_{j}=\emptyset$ for $i\neq j$.
+We say that $\mathcal{C}$ is trivial if either $k=1$, or $C_{i}$ is a singleton set for all $i$.
 We identify a community $C_{i}$ with the induced subgraph $G[C_{i}]=(C_{i},E(C_{i}))$ of the graph $G$ for every community $C_{i}$, where $E(C_{i}):=\{(v,w)\in E:v,w\in C_{i}\}$.
-The $E(\mathcal{C})=\cup_{i=1}^{k}E(C_{i})$ is the set of intra-community edges of the graph $G$ induced by the community detection $\mathcal{C}$.
+The $E(\mathcal{C})=\cup_{i=1}^{k}E(C_{i})$ is the set of intra-community edges induced by $\mathcal{C}$.
 The number of edges in $E(\mathcal{C})$ is denoted by $m(\mathcal{C})$.
-Also, $E(G)-E(C)=\{(u,v):u\in C_{i},\ v\in C_{j},\ i\neq j\}$ is the set of inter-community edges of $G$ induced by the community detection $\mathcal{C}$.
-[34].
+Also, $E(G)-E(C)=\{(u,v):u\in C_{i},\ v\in C_{j},\ i\neq j\}$ is the set of inter-community edges of $G$ induced by $\mathcal{C}$.
 The number of intra-community edges is more than the number of inter-community edges.
@@
 2.2 An introduction to quantum walks
-Let $G=(V(G),E(G))$ be a connected simple graph.
-We assign two opposite orientations on every undirected edge $(u,v)\in E(G)$.
-Therefore, from every undirected edge $(u,v)$ in $G$, we construct two directed edges $\overrightarrow{e}=\overrightarrow{(u,v)}$ and $\overrightarrow{e^{-1}}=\overrightarrow{(v,u)}$.
-The new directed graph is $\overrightarrow{G}$.
-Therefore, if $G$ has $m$ edges, then $\overrightarrow{G}$ has $2m$ directed edges.
-Let the edges $\overrightarrow{e_{1}},\overrightarrow{e_{2}},\dots,\overrightarrow{e_{m}},\overrightarrow{e_{1}^{-1}},\overrightarrow{e_{2}^{-1}},\dots,\overrightarrow{e_{m}^{-1}}$ are marked by the integers $0,1,2,\dots(2m-1)$.
-The construction of $\overrightarrow{G}$ from $G$ indicates that $d_{G}(v)=d^{(o)}_{\overrightarrow{G}}(v)=d^{(i)}_{\overrightarrow{G}}(v)$, where $d_{G}(v)$ is the degree of $v$ in $G$, as well as $d^{(o)}_{\overrightarrow{G}}(v),d^{(i)}_{\overrightarrow{G}}(v)$ are the out-degree and in-degree of $v$ in $\overrightarrow{G}$, respectively.
-Let a walker is moving on the directed graph $\overrightarrow{G}$ following the graph structure.
-At a particular time $t$, let the walker arrives at a vertex $v$.
-In the next step, the walker will move to one of the adjacent vertices of $v$.
-Hence, the walker must follow one of the outgoing edges $\overrightarrow{e}$ from $v$.
-Recall that, if out-degree of $v$ is $d_{v}^{(o)}$, then there are $d_{v}^{(o)}$ out-going edges from $v$.
-The walker will choose one of them with probability $p(\overrightarrow{e})$.
-Therefore, we can define a function $p:E(\overrightarrow{G})\rightarrow(0,1]$ satisfying
-$\sum_{\overrightarrow{e}:o(\overrightarrow{e})=v}p(\overrightarrow{e})=1,$ (1)
-for each vertex $v\in V$. If we assume that choosing an outgoing edge from $v$ is equiprobable, then $p(\overrightarrow{e})=\frac{1}{d^{(o)}_{v}}=\frac{1}{d_{v}}$, where $v=o(\overrightarrow{e})$.
-Now, we are in a position to define the Szegedy quantum walk.
-We represent the state of a walker by a quantum state vector.
-Corresponding to the directed graph $\overrightarrow{G}$ with $2m$ edges, we consider a Hilbert space $\mathcal{H}^{2m}$ of dimension $2m$.
-The basis of $\mathcal{H}^{2m}$ is $\left\{|0\rangle,|1\rangle,|2\rangle,\dots,|2m-1\rangle\right\}$, where
-[…]
+From a connected simple graph $G=(V(G),E(G))$ we construct a directed graph $\overrightarrow{G}$ by replacing every undirected edge $(u,v)$ by two directed edges $\overrightarrow{e}=\overrightarrow{(u,v)}$ and $\overrightarrow{e^{-1}}=\overrightarrow{(v,u)}$.
+Let the edges $\overrightarrow{e_{1}},\overrightarrow{e_{2}},\dots,\overrightarrow{e_{m}},\overrightarrow{e_{1}^{-1}},\overrightarrow{e_{2}^{-1}},\dots,\overrightarrow{e_{m}^{-1}}$ in $\overrightarrow{G}$ are marked by the integers $0,1,2,\dots(2m-1)$.
+The construction of $\overrightarrow{G}$ from $G$ indicates that $d_{G}(v)=d^{(o)}_{\overrightarrow{G}}(v)=d^{(i)}_{\overrightarrow{G}}(v)$.
+At time $t$, let the walker is at a vertex $v$.
+At time $(t+1)$, following one of the $d_{v}^{(o)}$ outgoing edges it reaches to a neighboring vertex.
+The walker chooses one of the out-going edges with probability $p(\overrightarrow{e})=\frac{1}{d^{(o)}_{v}}=\frac{1}{d_{v}}$, where $v=o(\overrightarrow{e})$.
+Therefore, $\sum_{\overrightarrow{e}:o(\overrightarrow{e})=v}p(\overrightarrow{e})=1$, for each vertex $v\in V$.
 […]
 $u_{\overrightarrow{e},\overrightarrow{f}}=\begin{cases}2\sqrt{p(\overrightarrow{e})p(\overrightarrow{f^{-1}})}-\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}&\text{if}~t(\overrightarrow{f})=o(\overrightarrow{e});\\ 0,&\text{otherwise}.\end{cases}$ (1)
-Here, $\delta_{(\overrightarrow{e},\overrightarrow{f})}$ is the Kronecker delta, which is defined by
-$\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}=\begin{cases}1&\text{if}~\overrightarrow{e}=\overrightarrow{f};\\ 0&\text{otherwise}.\end{cases}$ (4)
+Here, $\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}$ is the Kronecker delta, which is defined by
+$\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}=\begin{cases}1&\text{if}~\overrightarrow{e}=\overrightarrow{f^{-1}};\\ 0&\text{otherwise}.\end{cases}$ (2)
 […]
@@
 […]
-The procedure of clustering depends on the selection of the initial state.
-A different choice of initial state may lead us to another limiting probability distribution, which may not be useful for us.
-One such example is presented in Appendix A
-3 Community detection on graphs
-[…]
+The procedure of community detection depends on the selection of the initial state, which is further explored in Appendix A.
+At $t\textsuperscript{th}$ step of quantum walk for $t=0,1,2,\dots$, the state vector of the walker is
 […]
@@
 $\sum_{k=0}^{2m-1}|a_{t,k}|^{2}=1,~\text{for all}~t.$ (5)
+3 Community detection on graphs
 […]
 $p_{t}(\overrightarrow{k})=\braket{\psi^{\dagger}|P_{k}|\psi}=|\langle k|\psi\rangle|^{2}=|\alpha_{t,k}|^{2}.$ (6)
-[…]
-$\overrightarrow{P}_{t}=(p_{t}(\overrightarrow{e_{1}}),p_{t}(\overrightarrow{e_{2}}),\dots,p_{t}(\overrightarrow{e_{n}}),p_{t}(\overrightarrow{e_{1}^{-1}}),p_{t}(\overrightarrow{e_{2}^{-1}}),\dots,p_{t}(\overrightarrow{e_{m}^{-1}})).$ (9)
-Note that, every edge $e$ of the undirected graph $G$ contributed two directed edges $\overrightarrow{e}$ and $\overrightarrow{e^{-1}}$, in the directed graph $\overrightarrow{G}$.
-Thus, we can determine the edge probability $e$ in $G$ at $t$-th step of quantum walk as $p_{t}(e)=p_{t}(\overrightarrow{e})+p_{t}(\overrightarrow{e^{-1}})$.
-Therefore the edge probability vector of the undirected graph $G$ at the $t$-th step of quantum walk is represented by
+Using equation (6), we construct an edge probability vector for $\overrightarrow{G}$ at the $t$-th step of quantum walk as
+$\overrightarrow{P}_{t}$ $=\bigl(p_{t}(\overrightarrow{e}_{1}),\ldots,p_{t}(\overrightarrow{e}_{m}),p_{t}(\overrightarrow{e}_{1}^{-1}),\ldots,p_{t}(\overrightarrow{e}_{m}^{-1})\bigr).$ (7)
+Note that, every edge $e$ in $G$ contributes two directed edges $\overrightarrow{e}$ and $\overrightarrow{e^{-1}}$, in $\overrightarrow{G}$.
+Thus, we determine the probability of $e$ in $G$ at time $t$ as $p_{t}(e)=p_{t}(\overrightarrow{e})+p_{t}(\overrightarrow{e^{-1}})$.
+Therefore, the edge probability vector of $G$ at time $t$ as
 $P_{t}=(p_{t}(e_{1}),p_{t}(e_{2}),\dots,p_{t}(e_{m})),$ (8)
 where $p_{t}(e_{i})=p_{t}(\overrightarrow{e_{i}})+p_{t}(\overrightarrow{e_{i}^{-1}})$.
-Note that, due to the unitary dynamics, $P_{t}$ need not converge to a limiting distribution for $t\rightarrow\infty$ [36].
+Note that, due to the unitary dynamics, $P_{t}$ need not converge to a limiting distribution as $t\rightarrow\infty$ [42].
 But, the time average probability distribution converges.
@@
 Definition 1
-The edge probability vector at time $T$ of a graph $G$ is defined by
-$\begin{split}\Pi^{(T)}&=\left(\pi_{1}^{(T)},\pi_{2}^{(T)},\dots\pi_{m}^{(T)}\right)\\ &=\frac{1}{(T+1)}\sum_{t=0}^{T}P_{t}=\frac{1}{(T+1)}\left(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\dots,\sum_{t=0}^{T}p_{t}(e_{m})\right),\end{split}$
-where $\pi_{k}^{(T)}=\frac{1}{(T+1)}\sum_{t=0}^{T}p_{t}(e_{k})$ for $k=1,2,\dots m$.
-Note that, the edge probability vector at time $T$ changes over time. It proceeds to a limiting distribution when $T\rightarrow\infty$. Hence, we can define the edge probability vector of a graph.
+The edge-probability vector at time $T$ of a graph $G$ is defined by
+$\begin{split}&\Pi^{(T)}=\bigl(\pi_{1}^{(T)},\pi_{2}^{(T)},\ldots,\pi_{m}^{(T)}\bigr)=\frac{1}{T+1}\sum_{t=0}^{T}P_{t}\\ &=\frac{1}{T+1}\bigl(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\ldots,\sum_{t=0}^{T}p_{t}(e_{m})\bigr),\end{split}$
+where $\pi_{k}^{(T)}=\frac{1}{T+1}\sum_{t=0}^{T}p_{t}(e_{k})$ for $k=1,2,\ldots,m$.
+Note that, the edge probability vector proceeds to a limiting distribution when $T\rightarrow\infty$, which we define as the edge probability vector of a graph.
 Definition 2
-The edge-probability vector of a graph $G$ is defined by
-$\Pi=(p(e_{1}),p(e_{2}),\dots,p(e_{m}))=\lim_{T\to\infty}\frac{1}{(T+1)}\sum_{t=0}^{T}P_{t}=\lim_{T\to\infty}\frac{1}{(T+1)}\left(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\dots,\sum_{t=0}^{T}p_{t}(e_{m})\right).$
-For our computation, rather than calculating $T\to\infty$, we find $\frac{1}{T+1}\sum_{t=0}^{T}P_{t}$ for a finite $T$ to construct the limiting probability distribution.
-We apply the standard Euclidean distance to differentiate two edges probability vectors $\Pi^{(T)}$ and $\Pi^{(T-1)}$, which is determined by
-$\begin{split}\left\lVert\Pi^{(T)}-\Pi^{(T-1)}\right\rVert&=\left\lVert\frac{1}{T+1}\sum_{t=0}^{T}P_{t}-\frac{1}{T}\sum_{t=0}^{T-1}P_{t}\right\rVert\\[6.0pt] &=\sqrt{(\pi_{1}^{(T)}-\pi_{1}^{(T-1)})^{2}+(\pi_{2}^{(T)}-\pi_{2}^{(T-1)})^{2}+\cdots+(\pi_{m}^{(T)}-\pi_{m}^{(T-1)})^{2}}.\end{split}$ (11)
-To construct the edge probability vector of a graph $G$, we compute a sequence of edge probability vectors $\{\Pi^{(t)}\}_{t=0}^{T}$ for time $t$, where
+For a graph $G$ the edge-probability vector
+$\begin{split}&\Pi=\bigl(p(e_{1}),p(e_{2}),\ldots,p(e_{m})\bigr)=\lim_{T\to\infty}\frac{1}{T+1}\sum_{t=0}^{T}P_{t}\\ &=\lim_{T\to\infty}\frac{1}{T+1}\Bigl(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\ldots,\sum_{t=0}^{T}p_{t}(e_{m})\Bigr).\end{split}$
+For our computation, rather than calculating $T\to\infty$, we find $\frac{1}{T+1}\sum_{t=0}^{T}P_{t}$ for a finite $T$.
+We apply the standard Euclidean distance to differentiate two edge probability vectors $\Pi^{(T)}$ and $\Pi^{(T-1)}$, which is
+$\left\lVert\Pi^{(T)}-\Pi^{(T-1)}\right\rVert$ $=\left\lVert\frac{1}{T+1}\sum_{t=0}^{T}P_{t}-\frac{1}{T}\sum_{t=0}^{T-1}P_{t}\right\rVert.$ (9)
+To work out the edge probability vector of $G$, we compute a sequence of edge probability vectors $\{\Pi^{(t)}\}_{t=0}^{T}$, where
 $\left|\Pi^{(T)}-\Pi^{(T-1)}\right|<\epsilon.$ (10)
-Here, we consider $\epsilon$ as a numerical error which is usually a very small number.
-For our numerical experiments, we consider $\epsilon=0.0001$.
-The consideration of numerical error assists us to find the accurate value of $\Pi$ to three decimal places.
+In our numerical experiments, $\epsilon=0.0001$, which assists us to find the accurate value of $\Pi$ to three decimal places.
 In 1(b), we present the edge-probabilities of the graph in 1(a).
-The convergence of the probability vectors are discussed in Appendix B.
+The convergence of the probability vectors are discussed in Appendix.
 (a) Graph $G$
@@
 Figure 1: Different steps for generating communities on the graph $G$, depicted in Sub-figure 1(a), are presented.
-Now, we include a number graph theoretic ideas in construction of our communities.
+Now, we include a few graph-theoretic ideas.
 A vertex with degree $1$ is adjacent to only one vertex, which may belong to a community.
@@
 But, a vertex with greater degree is adjacent to many other vertices.
-Thus, it belong to only one community.
+Thus, it belongs to only one community.
 It insists us to arrange the vertices in a descending order of their degrees.
-In every step, we consider the highest degree vertex among the vertices which are not allotted to a community and generate a community for it.
-In addition, the above paragraph justifies that the community containing a vertex with highest degree also contains all its neighboring vertices.
-But, there may be a number of vertices in a community, which are not an immediate neighbor of the vertex with highest degree in it.
+In every step, we consider the highest degree vertex among the vertices which are not allotted to a community.
+There may be a number of vertices in a community, which are not an immediate neighbor of the vertex with highest degree in it.
 Also, there are non-neighboring vertices which are not in the community.
@@
 Therefore, we need a condition for including a given vertex in a particular community.
-Recall from subsection 2.1 that $N(v_{i})=\{v:(u,v)\in E(G)\}$.
+Recall that $N(v_{i})=\{v:(u,v)\in E(G)\}$.
 Let $\operatorname{Nbd}(v_{i})$ be the neighborhood of $v_{i}$ in the community of vertices containing $v_{i}$.
 If $|N(v_{j})\cap\operatorname{Nbd}(v_{i})|\geq\frac{|N(v_{j})|}{2}$ holds, we assume that $v_{j}$ is in the community of $v_{i}$.
-We make the final selection of a vertex in a community based on the edge probability vector $\Pi$ generated by quantum walk.
+The final selection of a vertex in a community depends on the edge probability vector $\Pi$.
 We determine the shortest path between $v_{i}$ and $v_{j}$.
-A path consists of a number of incident edges.
-The walker follows the path means it follows all the subsequent edges.
+The walker follows the path means it follows all the consecutive edges.
 As there is no non-Markovian componenet in our quantum walk, the probability of getting the walker on the edges of a path is independent to each other.
-Therefore, the probability of getting the walker on a path is the product of the probability of getting the walker on the edges.
+Therefore, the probability of getting the walker on a path is the product of the probabilities.
 It leads us to define the path weight between two vertices.
@@
 where $p(e_{i})$ is the probability corresponding to the edge $e_{i}$ in the edge probability vector $\Pi$.
-Note that, as all $0<p(e_{i})<1$ and $d_{G}(v)\geq 1$ we have $0<PW(u,v)<1$ for all pairs of vertices $u$ and $v$.
-The explanation for dividing $\frac{1}{d_{G}(v)}$ will be discussed after Procedure Procedure 1.
-Now, we are in a position to determine the communities in a network.
-Given a vertex $v_{i}$ the community consists of $v_{i}$ is denoted by $C(v_{i})$.
+Let $v_{i}$ be a vertex of maximum degree among the vertices which are not included in a community.
+The community consists of $v_{i}$ is denoted by $C(v_{i})$.
 Initially, $C(v_{i})=\{v_{i}\}$.
-Then, $C(v_{i})$ grows maintaining some conditions.
-We choose the vertices $v_{i}$ according to their degrees.
-Our method is described in the procedure mentioned below.
+The Procedure mentioned below add the neighboring vertices of $v_{i}$ in $C(v_{i})$.
 Procedure 1
-Recall that $N(v)$ denotes the subset of vertices which are adjacent to $v$ in $G$. The procedure for detecting the communities consists of the following steps:
+Follow the steps mentioned below:
   1. 1.
-    Arrange the vertices of the graph $G$ in the descending order of the vertex degrees.
-Therefore, the vertex set of $G$ is $V=\{v_{1},v_{2},\dots v_{n}\}$, where $d_{G}(v_{i})\geq d_{G}(v_{i+1})$ for $i=1,2,\dots(n-1)$.
+    Arrange the vertices of $G$ as $V=\{v_{1},v_{2},\dots v_{n}\}$, where $d_{G}(v_{i})\geq d_{G}(v_{i+1})$.
   2. 2.
-    Repeat the following steps if $V\neq\emptyset$.
+    Repeat the following steps if $V\neq\emptyset$:
     1. (a)
@@
     2. (b)
-      Construct the set of vertices $\operatorname{Nbd}(v_{i})=N(v_{i})$.
+      Construct $\operatorname{Nbd}(v_{i})=N(v_{i})$.
     3. (c)
@@
     6. (f)
-      Update the set $V$ by removing all the vertices in $C(v_{i})$.
+      Update the set $V$ by $V-C(v_{i})$.
   3. 3.
-    Return the communities of $C(v_{i})$ which are formed in Step 2.
-It is clear from the Procedure 1 mentioned above that we construct a set of vertices $\operatorname{Nbd}(v_{i})$ based on equation (14), which assists us to construct the community $C(v_{i})$.
-The quantum walk is useful to calculate the probability values in finding $PW(v_{i},v_{j})$ mentioned in equation (13).
-In the above procedure $q$ is a parameter.
-If $q=0$ then the number of communities is equal to the number of vertices in the graph, that is every community consists of single vertices.
+    Return the communities of $C(v_{i})$.
+In Procedure 1 the quantum walk is useful to calculate the probability values in finding $PW(v_{i},v_{j})$ mentioned in Definition 3.
+Also, $q$ is a parameter.
+If $q=0$ then the number of communities is equal to the number of vertices in the graph.
 When $q>\max_{(u,v)\in E(G)}\{PW(u,v)\}$ then $C(v_{i})=\operatorname{Nbd}(v_{i})$.
-In equation (13), we multiply $\frac{1}{d_{G}(v)}$ to normalize node influence.
+In Definition 3, we multiply $\frac{1}{d_{G}(v)}$ to normalize node influence.
 It provides more preference to a high-degree node to be added in $C(v_{i})$, because a node with higher degree reduces the product $\prod_{i=1}^{n}p(e_{i})$ than a vertex with lower degree.
@@
     For any vertex $v_{j}\in C(v_{i})$ if $d_{C(v_{i})}(v_{j})\leq\frac{d_{G}(v_{j})}{2}$ then remove $v_{j}$ from $C(v_{i})$.
-Here, $C(v_{i})$ denotes the induced subgraph of $G$, generated by $C(v_{i})$.
 Therefore, our method of community detection is a combination of the following steps, mentioned in the below procedure.
 Procedure 3
-Let $G$ be a simple undirected graph with $n$ vertices and $m$ edges. To find the communities in $G$ perform the following steps:
+To find the communities in a simple graph $G$ perform the following steps:
   1. 1.
-    Generation of a directed graph: Convert the graph $G$ to a directed graph $\overrightarrow{G}$ by assigning opposite orientations on every undirected edge.
+    Generation of a directed graph: Generate a directed graph $\overrightarrow{G}$ from $G$ by assigning opposite orientations on every undirected edge.
   2. 2.
@@
   4. 4.
-    Generation of limiting probability distribution: For a fixed error term $\epsilon$, we apply equation (12) to stop quantum walk and generate the limiting probability distribution $\Pi$ on the edges mentioned in Definition 2.
+    Generation of limiting probability distribution: For a fixed error term $\epsilon$, we apply equation (10) to stop quantum walk and generate the limiting probability distribution $\Pi$ mentioned in Definition 2.
   5. 5.
@@
 4 Numerical experiments
-We have already mentioned that not all graphs admits community structure in it.
-We consider a number of graphs with a community structure.
-We also consider a graph having no trivial community structure.
-We initiate our discussion with a few standard examples.
-Some of the graphs considered in this article are the random graphs.
-Due to their generation procedure, they may have directed edges or loops.
+We consider a number of standard examples of graphs with a community structure, which are considered in literature.
+Some of them are the random graphs, which may have directed edges or loops.
 But, our procedure is applicable for undirected simple graphs.
-During the network generation process, we remove the loops and directed edges if exist.
-For every graph we include a figure with at least four sub-figures. Below, we describe them:
-  1. 1.
-    In Sub-figure (a), we draw the example graph, which we consider as our input.
-  2. 2.
-    In Sub-figure (b), we include a bar diagram representing the limiting edge-probability distribution $\Pi$, mentioned in Definition 2 after completion of quantum walk.
-  3. 3.
-    In Sub-figure (c), we draw a bar diagram mentioning the path weights between selected vertices in the $\operatorname{Nbd}$ of the community generating vertex.
-The equation of path probability is mentioned in equation (13).
-  4. 4.
-    In Sub-figure (d), we draw the graph with communities in it. The vertices in same community have same color. There are communities with single vertices. We mark all of them in gray.
-We use NetworkX, a Python library, for all the calculations related to graphs [37].
+Hence, we remove the loops and directed edges if exist.
+We use NetworkX, a Python library, for all the computations related to graphs [43].
 4.1 Barbell graphs
-A barbell graph [38] is generated by joining $r$ copies of a complete graph $K_{s}$ with a particular vertex, called central vertex.
-We join the central vertex to the $r$ vertices belonging to different $K_{s}$ by $r$ edges.
-For example in 2(a), we depict a barbell graph, that has two complete graphs each of five vertices.
-We add two vertices from each of the $K_{5}$ graphs with a vertex $5$ by two edges.
-Here vertex $5$ acts as the central vertex.
-Therefore, the barbell graph under our consideration has eleven vertices and twenty-two edges.
-Considering $q=\frac{1}{m}=0.0455$ in Procedure 3, we obtain three distinct communities in the barbel graph network.
-Two distinct communities arise from two complete graphs.The central vertex forms a community with a single vertex.
+In 2(a), we consider a barbell graph [44] which is generated by joining two copies of the complete graph $K_{5}$ with a central vertex by two edges.
+This graph has eleven vertices and twenty-two edges.
+Considering $q=\frac{1}{m}=0.0455$ in Procedure 3, we obtain three distinct communities in the barbel graph.
+Two distinct communities arise from two complete graphs.
+The central vertex forms a community with a single vertex.
 The coverage values corresponding to $q=0.0455$ is $0.909$.
-In 2(d) we present the communities generated by Procedure 3.
+In 2(c), we present the communities generated by Procedure 3.
 (a) A barbell graph where two complete graph $k_{5}$ are joined to a central vertex $5$ by two edge $(5,6),(5,4)$ .
-(b) Edge probabilities in the limiting distribution $\Pi$ are depicted as bar diagram for the graph in Subfigure 2(a).
-(c) Path weight between different vertices of the for generating the communities from the graph in 2(a).
-(d) Communities in the graph of Subfigure 2(a) are extracted with three different colors.
+(b) Path weight between different vertices of the for generating the communities from the graph in 2(a).
+(c) Communities in the graph of Subfigure 2(a) are extracted with three different colors.
 Figure 2: Community detection on a barbell graph
 4.2 Relaxed caveman graph
-The idea of caveman graphs arises from the social networks [39, 40].
-The connected caveman graph is generated by creating $k$ complete subgraphs with equal number of vertices in each.
-Then a single edge in each clique is rewired to a node in an adjacent clique.
-These cliques are called caves.
-The relaxed caveman graph is a variant of the connected caveman graphs.
-The network is composed of $k$ nodes organized into $k$ caves.
-For every edge there is a probability $p$ of rewiring, that is removing an edge and adding it to any vertex keeping one end point fixed.
-A relaxed caveman graph is depicted in 3(a).
-The example graph, depicted in 3(a), has $15$ vertices and $30$ edges.
-The vertices are distributed into $3$ caves.
-Every cave contains $5$ vertices in it.
-We consider that the rewiring probability is $0.2$.
-The example graph has $4$ inter-community edges.
-Note that, in this case $\frac{1}{m}=0.033$.
-The communities are marked with three different colors in Sub-figure 3.
+The connected caveman graph [45, 46] is generated by joining complete subgraphs each with $k$ vertices.
+An edge in a clique is rewired to a node in an adjacent clique with probability $p$.
+In 3(a) we depict a relaxed caveman graph having $30$ edges, and $15$ vertices which are equally distributed in $3$ caves.
+We assume $p=0.2$ which creates $4$ inter-community edges.
+Putting $q=\frac{1}{m}=0.033$, we find three communities which are presented in Sub-figure 3.
 Decreasing the value of $q_{2}$ to $0.0166$ we find more communities in our network.
-Coverage are $0.866$ and $0.766$ for $q=0.0333$ and $0.0166$ respectively.
+Coverage are $0.866$ and $0.766$ for $q=0.0333$ and $0.0166$, respectively.
 (a) Relaxed caveman graph having three caves with rewiring probability $p=0.2$.
@@
 4.3 Planted $l$-partition graph
-The planted partition models of networks offer another controllable artificial input for evaluating and benchmark the community detection algorithms.
-The Erdős-Rényi model of random graphs [41] with $n$ vertices is generated by choosing each of the $\binom{n}{2}$ possible edges independently and randomly to be included in the graph with probability $p$.
-The planted $l$-partition model is a modified version of the Erdős–Rényi model, which is specifically designed to generate communities in graph.
-To generate $l$ communities in a graph with $k$ vertices in each, we need $n=l\times k$ vertices.
-Now, we fix two probabilities $p$ and $q$ with $p>q$.
-Existence of an edge between two vertices belonging to same cluster is determined by probability $p$.
-An edge between two vertices from two different cluster occurs in probability $q$.
-A planted $l$-partition graph is depicted in 4(a).
-We consider a sample of a planted partition graph, which in 4(a).
-It has twelve vertices and fourteen edges.
-Also, values of $p_{1}$ and $p_{2}$ are $0.4$ and $0.1$, respectively.
-The vertices are divided into three subset each having four vertices in it.
-The vertices with the highest degrees in the graph are identified as 0, and $7$.
-Among them, vertex 0 is randomly selected as the first vertex to initiate Procedure 1 for constructing a community.
-A bar diagram is employed to illustrate the route probability distribution.
-Based on the graphs structural connectivity, a distinct community can be obtained for $q>0.071$.
+The planted $l$-partition graph [5, 47] has $n=l\times k$ vertices distributed into $l$ communities with $k$ vertices in each.
+Let $p$ and $q$ with $p>q$ be the probability of existence of the intra-community and inter-community edges, respectively.
+4(a) contains a planted $l$-partition graph with $14$ edges and $12$ vertices which are distributed into $3$ communities each having $4$ vertices in it.
+Also, $p_{1}=0.4$ and $p_{2}=0.1$, respectively.
+One of the highest degree vertex 0 is randomly selected as the first vertex to initiate Procedure 1.
+Considering $q>0.071$, we obtain three communities, where a community contain a single vertex.
 The coverage values corresponding to $q=0.071$ is $0.857$.
 (a) A planted partition graph with $p_{1}=0.4$ and $p_{2}=0.1$.
-(b) Edge probability in the limiting distribution $\Pi$ is depicted as bar diagram for the graph in Subfigure 4(a).
-(c) Path weight between different vertices of the for generating the communities from the graph in 4(a).
-(d) Communities in the graph of Subfigure 4(a) are extracted with three different colors.
+(b) Path weight between different vertices of the for generating the communities from the graph in 4(a).
+(c) Communities in the graph of Subfigure 4(a) are extracted with three different colors.
 Figure 4: Community detection in a planted partition graph.
@@
 The Zachary’s karate-club network [48] is another well-studied graph considered in the literature of community detection [40, 39].
-This is a social network of university karate club members.
-Each of the $34$ vertices represents a club member.
-There are $78$ edges represent the relations between the members.
-Article [33] presents three communities in the Zachary’s karate-club network.
-In contrast, article [32] utilize a different community detection technique to analyze the same network that found two communities only.
-Considering $q=\frac{1}{m}=0.0128$ in Procedure 3, we obtain three distinct communities in the Zachary’s karate-club network.
+It has $34$ vertices and $78$ edges.
+Article [40] presents three communities in it.
+In contrast, article [39] consider the same network that found two communities only.
+Considering $q=\frac{1}{m}=0.0128$ in Procedure 3, we obtain three communities in this network.
 Two communities have fourteen and eighteen vertices, and one community consists of a single vertex.
-The coverage values corresponding to $q=0.0128$ is found to be $0.859$.
-In 5(d) we present the communities determined by Procedure 3.
-In contrast, the article [33] determines three communities having five, ten, and nineteen vertices.
-Furthermore, article [32] develops two communities consist of sixteen and eighteen vertices.
+The coverage values corresponding to $q=0.0128$ is $0.859$.
+In 5(c), we present the communities determined by Procedure 3.
 (a) Zachary’s karate club network.
-(b) The bar diagram for the graph in Subfigure 5(a) represents the edge probability in the limiting distribution $\Pi$.
-(c) In 5(a), the path weight between various vertices of the graph is used to generate the communities.
-(d) Three distinct colors are used to extract communities from the graph in Subfigure 5(a).
+(b) In 5(a), the path weight between various vertices of the graph is used to generate the communities.
+(c) Three distinct colors are used to extract communities from the graph in Subfigure 5(a).
 Figure 5: Karate club graph are considered to generate communities.
 4.5 Les Misérable graph
-6(a) represents the “Les Misérable” graph.
-This network was demonstrated in [43] considering the characters in Victor Hugo’s novel “Les Misérable” as vertices.
-Any pair of characters that occur in the same chapter of the novel are connected by an edge.
-This graph has seventy seven vertices and two hundred fifty four edges.
-Article [33] considers the Les Misérable graph for community detection.
-It determines six communities in the graph.
-We use Procedure 3 on the Les Misérable graph considering $q=\frac{1}{m}=0.0039$ for determining the communities.
-It generates eighteen communities in the graph.
-Among them there are three communities with vertices eight, twelve, and forty-four.
-Remaining fifteen communities have single vertices in each.
-In 6(d), we present the communities in the Les Misérable graph determined by our procedure.
-The coverage values corresponding to $q=0.0039$ is found to be $0.752$.
+6(a) represents the “Les Misérable” graph [49]. having $77$ vertices and $254$ edges.
+Article [40] determines six communities in the graph.
+We use Procedure 3 considering $q=\frac{1}{m}=0.0039$ for determining the communities, that generates eighteen communities.
+Among them there are three communities with $8$, $12$, and $44$ vertices.
+Remaining $15$ communities have single vertices in each.
+In 6(c), we present these communities.
+The coverage values corresponding to $q=0.0039$ is $0.752$.
 (a) Les Misérable graph
-(b) The graph in Subfigure 6(a) has a bar diagram that shows the edge probability in the limiting distribution $\Pi$.
-(c) In 6(a), the path weight between various vertices of the graph is used to generate the communities.
-(d) The three communities are represented by three distinct colors in the graph shown in Subfigure 6(a), while all other single-node clusters are shown by gray.
+(b) In 6(a), the path weight between various vertices of the graph is used to generate the communities.
+(c) The three communities are represented by three distinct colors in the graph shown in Subfigure 6(a), while all other single-node clusters are shown by gray.
 Figure 6: Les Misérable graph are utilized to generate communities.
 4.6 Dolphin’s social network
-7(a) illustrates the dolphin’s social network which represents a social network among sixty two bottlenose dolphins, which are considered as nodes.
-Each of the one hundred fifty nine edges signifies a frequent social interaction between two dolphins spotted together more often than by chance [44].
-Article [45] determines four communities in the dolphin’s social network.
-These communities consists of six, thirteen, twenty two and twenty one vertices.
-Considering $q=\frac{1}{m}=0.0063$ in Procedure 3, we obtain five distinct communities in dolphin social network.
-There are three communities having twelve, twenty-one, and twenty-seven vertices, and two communities consists of a single vertex in each.
+7(a) illustrates the dolphin’s social network [50]. Article [51] determines four communities in the network, that consists of $6$, $13$, $22$, and $21$ vertices.
+Considering $q=\frac{1}{m}=0.0063$ in Procedure 3, we obtain $5$ communities.
+There are three communities having $12$, 2121, and $27$ vertices, and two communities consists of a single vertex in each.
 These single-vertex communities exist because the respective vertices are associated with several communities and hence cannot be unequivocally allocated to any one of the three communities.
@@
 (a) Dolphin social network
-(b) The bar diagram for the graph in Subfigure 7(a) represents the edge probability in the limiting distribution $\Pi$.
-(c) In 7(a), the path weight between various vertices of the graph is used to generate the communities.
-(d) The three communities are represented by three distinct colors in the graph shown in Subfigure 7(a), while the other two single-node communities are shown by gray.
+(b) In 7(a), the path weight between various vertices of the graph is used to generate the communities.
+(c) The three communities are represented by three distinct colors in the graph shown in Subfigure 7(a), while the other two single-node communities are shown by gray.
 Figure 7: Dolphin Network are utilized to generate communities.
-4.7 Euroroad network
-We consider Euroroad data from [46] for our experiment.
-The road interactions are represented by the vertices in the network.
-Every edge denotes a road that links two junctions.
-Our road network contains $1174$ nodes and $1417$ edges, which is not connected.
-Hence, we study the largest connected component in this network containing $1039$ vertices and $1305$ edges.
-This component is depicted in 8(a).
-Applying Procedure 3, we find three hundred eighty-nine communities. The largest community consists of thirty-five vertices. The coverage values corresponding to $q=0.00077$ is found to be $0.540$.
-(a) Connected Road network
-(b) The bar diagram for the graph in Subfigure 8(a) represents the edge probability in the limiting distribution $\Pi$.
-(c) In 8(a), the path weight between various vertices of the graph is used to generate the communities.
-(d) All communities are represented by distinct colors in the graph shown in Subfigure 8(a), while the single-node communities are shown by gray.
-Figure 8: Road Network are utilized to generate communities.
 5 Conclusion
@@
 There are multiple models of quantum walks in the literature.
-We employ the Szegedy quantum walks on graphs for our investigation, as it is defined on all graphs.
+We employ the Szegedy quantum walks on graphs for our investigation.
 The limiting probability distribution of the Szegedy quantum walks assists us in finding the communities.
@@
 We assume that the walker begins its journey from the vertices with the highest degree.
-Therefore, the initial quantum state as well as the limiting probability distribution depends on the structure of the graph.
-The limiting probability distribution provides the edge probabilities for each edge in the graph.
+The initial quantum state as well as the limiting probability distribution depends on the structure of the graph.
 We use it in Procedure 3 to extract the communities from the graph.
@@
 We apply Procedure 3 on different graphs to identify the communities in them.
-The problem of community detection is an important problem in the study of networks.
-The communities assists us to understand the structure of the networks.
-In the future, the interested reader may apply our community detection procedure in the investigation of different complex networks.
+In the theory of complex network, the problem of community detection is an important problem, because the communities assists us to understand the structure of the networks.
+In the future, the interested reader may apply our community detection procedure for investigating the structure of different complex networks.
 Acknowledgment
@@
 0123456
-(a) An example graph
-[…]
+(a) A barbell graph.
+Edge $\Pi^{\prime}$ $\Pi$
 (0,1) 0.1274 0.1625
@@
 (5,6) 0.1274 0.1116
-(b) Limiting probability distributions on the graph in 9(a) for different choice of initial state.
-Figure 9: We calculate limiting probability distributions mentioned in Definition 2 for a graph in 9(a) with different choice of initial states, which are mentioned in equation (5) and (15).
-We receive two different probability distributions in two different cases.
-They are in the table 9(b).
-In this article, we consider a particular initial state, which is mentioned in equation (5).
-But, we had many other options of initial state of the walker.
-For example the equal superposition state as follows:
+(b) Limiting distributions.
+Figure 8: A barbell graph and different limiting probability distributions for different initial states.
+We consider an initial state, mentioned in equation (3). Consider another state
 […]
-Equation (15) indicates that the walker will start from all the edges selected uniformly.
 […]
-Appendix B Convergence of probability distributions
-In Definition 2, we claim that the edge probability vectors, mentioned in Definition 1 will converge to a vector over time.
-The Markovian nature of the considered quantum walk is the key reason for this convergence.
-We present this incident via a number of diagrams.
-depicted in Figure 10.
-We consider six network for our investigations.
-From every graph we select a random sample of four edges.
-For different time instance every edge has an edge probability value.
-We observe that all these probability values converges over time for any particular edge.
-Find Figure 10 for the detail descriptions.
+The edge probability vectors, mentioned in Definition 1, converges to a limiting distribution, mentioned in Definition 2, which we can observe numerically.
+We consider a random sample of $4$ edges $\{(1,12),(23,24),(3,4),(6,7)\}$ from the karate club graph, which is depicted in 5(a).
+As none of them is incident to the highest degree vertices their probability values at $t=0$ is 0.
+Upto $t=50$ we present the changes in probability values which is depicted in Figure 9.
 (a) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the barbell graph, depicted in 2(a). We observe the convergence from $T=743$.
@@
 (f) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the dolphins Graph, depicted in 7(a). We observe the convergence from $T=705$.
-(g) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the road network, depicted in 8(a). We observe the convergence from $T=322$.
 Figure 9: Convergence of edge probability distributions, which is mentioned in Definition 2.